Circular No. 6854

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980329 J. in ’t Zand and J. Heise, Space Research Organization of the Nether-

lands, Utrecht; L. Piro, Istituto Astrofisica Spaziale, CNR, Rome; L. A. Antonelli, BeppoSAX Science Data Center, Rome; M. Daniele, G. Gennaro, G. Spoliti, and V. Torroni, BeppoSAX Science Operation Center, Rome; and E. Pian, Istituto di Tecnologie e Studio delle Radiazioni Extraterrestri, CNR, Bologna, report: “The BeppoSAX Wide Field Camera error box of GRB 980329 (IAUC 6853) was observed with the BeppoSAX Narrow Field Instruments (NFI) starting on Mar. 29.449 UT (7.0 hr after the burst). A previously unknown x-ray source, 1SAX J0702.6+3850, was detected by the MECS (units 2 and 3) at α = 7^{h}02^{m}36^{s}, δ = +38^{o}50′.4 (equinox 2000.0; preliminary error radius 1′). This position is 1′.0 from the centroid of the WFC position. The source 1SAX J0702.6+3850 was brightest at the start of the observation; in the first five hours, the average countrate was 0.013 ± 0.0015 count/s (1.6–10 keV) in the two MECS units, corresponding to (7.8 ± 0.9) × 10^{−13} erg cm^{−2} s^{−1}. The source faded by a factor of about 3 over 14 hr. We propose 1SAX J0702.6+3850 as the x-ray afterglow of GRB 980329.”

AX J0051–73.3 = RX J0050.7–7316 P. Kahabka, University of Amsterdam, communicates: “The new x-ray

pulsar AX J0051–73.3 in the Small Magellanic Cloud reported on IAUC 6853 coincides in position within the 1′ error circle with the ROSAT source RX J0050.7–7316. The source was detected with the PSPC on 1992 Apr. 15–24 and with the HRI on 1995 May 18–June 5. From the HRI observation, we derive a position of α = 0^{h}50^{m}44^{s}.5, δ = −73^{o}16′03′′ (equinox 2000.0; ± 4′′ at 90-percent confidence).”

COMET C/1995 O1 (HALE-BOPP) Visual m_{1} estimates: Feb. 24.02 UT, 8.4 (M. Lehky, Puerto Escondido,

Falcon, Venezuela, 0.07-m refractor); Mar. 3.01, 8.4 (R. J. Bouma, Merida, Venezuela, 15×80 binoculars); 15.25, 8.8 (N. Biver, Honolulu, HI, 0.26m reflector); 19.44, 7.9 (D. A. J. Seargent, The Entrance, N.S.W., 10×50 binoculars); 25.49, 8.9 (M. Mattiazzo, Wallaroo, S. Australia, 0.20-m reflector); 26.99, 9.1 (J. G. de S. Aguiar, Campinas, Brazil, 0.23-m reflector).

1998 March 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green